Exercise 20.6 (Algorithmic)
Ordering and Carrying Costs

Ottis, Inc., uses 683,400 plastic housing units each year in its production of paper shredders. The cost of placing an order is $33. The cost of holding one unit of inventory for one year is $12. Currently, Ottis places 170 orders of 4,020 plastic housing units per year.

Required:

1. Compute the annual ordering cost.

$

2. Compute the annual carrying cost.

$

3. Compute the cost of Ottis’s current inventory policy.

$

Is this the minimum cost?

- Select your answer -YesNo
Answer #1
Annual ordering cost = (683400/4020)*33
Annual ordering cost = 170*33
Annual ordering cost = $ 5,610/.
Annual carrying cost = (4020/2)*12
Annual carrying cost = $ 24,120/.
Total cost of Ottis’s current inventory policy = (5610+24120)
Total cost of Ottis’s current inventory policy = $ 29,730/.
This is the minimum cost, Yes